How do you find a quadratic function to model the values in the table? Predict the value of y for x=6.
x y
-1 -2
0 4
3 -26
Could you also help me with finding a quadratic model for th set of values.
(-2,19)(0,3)(4,19) Thanks alot. Forgot all these things that we did at the beginning of the year.

Answer by Nate(3500)

(-2,19)(0,3)(4,19) with vertex (1,f(1))
f(x) = a(x - h)^2 + k is the standard form
f(x) = a(x - 1)^2 + f(1)
3 = a(-1)^2 + f(1) PLUG
3 - f(1) = a
~ Equation ~
f(x) = (3 - f(1))(x - 1)^2 + f(1)
19 = (3 - f(1))(-3)^2 + f(1) PLUG
19 = 27 - 9f(1) + f(1)
-8 = -8f(1)
1 = f(1)
f(x) = 2(x - 1)^2 + 1
~~~ x = 6 ... y or f(x) = ?
f(6) = 2(5)^2 + 1 = 50 + 1 = 51